OpenStudy (anonymous):

How do you find 3tan^3x=tanx?

OpenStudy (anonymous):

Simplify both sides with tanx 3tan^2x=1 , tan^2x=1/3 and root both side , \[\left| tanx \right|=1/\sqrt{3}\] So go on from here

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No, don't do that. Unless you want to lose half your solutions.

OpenStudy (anonymous):

\[3\tan^3x = \tan x \iff \tan x\ (3\tan^2x -1) = 0\] \[\implies \tan x = 0 \implies x = ...\] OR \[\implies \tan^2x = \frac{1}{3} \implies \tan x = \pm \sqrt{\frac{1}{3}}\]